OpenStudy (anonymous):

Hello I need help with this equation simplify (2 rad.8) (3 rad.6) + 19 rad.48

OpenStudy (anonymous):

23997878253756106444997335735942416102218125380085922630175553 :)

OpenStudy (anonymous):

what does rad.8 mean

OpenStudy (anonymous):

to the power of 8

OpenStudy (anonymous):

Okay (\[(2\sqrt{8}) (2\sqrt{6}) + 19 \sqrt{48}\]

OpenStudy (anonymous):

3 \[sorry (3\sqrt{6 })\]

OpenStudy (anonymous):

46\[\sqrt{12}\]

OpenStudy (phi):

for the first part, you can change the order of the multiplies \( 2 \cdot 3 \sqrt{8}\cdot \sqrt{6} \) and use the rule \[ \sqrt{a} \sqrt{b} = \sqrt{a b} \]

OpenStudy (anonymous):

should i simplify

OpenStudy (anonymous):

its 46\[\sqrt{12}\]

OpenStudy (anonymous):

or 92\[\sqrt{3}\]

OpenStudy (anonymous):

if there are any options i can choose

OpenStudy (anonymous):

so would it be like \[2 \sqrt{2}\sqrt{4} and 3 \sqrt{2}\sqrt{3} + 19 \sqrt{2}\sqrt{14}\]

OpenStudy (anonymous):

ya likely

OpenStudy (anonymous):

sorry rathan reddy the answers are 1) 25 rad.3 2) 19 rad. 6 3) 19 rad 12 + 2rad 48 4) 100 rad 3 5) 100 rad 12 + 3 rad 48

OpenStudy (phi):

yes you could do that (but the "and" is really multiply) but I was thinking just write the first part as \[ 2 \cdot 3 \sqrt{8}\cdot \sqrt{6} = 6 \sqrt{48}\] so that the problem is \[ 6 \sqrt{48} + 19 \sqrt{48} \] you have 6 square roots (of 48) and add 19 more, how many do you have ?

OpenStudy (phi):

you should get 25 sqrt(48) now simplify sqrt(48) can you do that ?

OpenStudy (anonymous):

for 48 I got 1*48 2*14 4*12 6*8

OpenStudy (anonymous):

i think its option 3.

OpenStudy (phi):

that way I simplify square roots is break them into "prime factors" and look for PAIRs of numbers: I start by dividing by little numbers: 48 = 2* 24 = 2* 2* 12 = 2*2*2*6 = 2*2*2*2*3 we have two pairs of twos: (2*2) * (2*2) * 3 pull each pair outside the square root and put just one of the pair on the outside. you get \[ \sqrt{ 2\cdot 2\cdot 2\cdot 2\cdot 3} = 2\cdot 2\sqrt{3}\]

OpenStudy (phi):

which is 4 sqrt(3) so you have 25 * 4 * sqrt(3) which simplifies

OpenStudy (anonymous):

sorry its 100 rad.3 i will simplyfy it now

OpenStudy (anonymous):

I don't understand the pull each pair outside the square root and put just one of the pair on the outside. you get 2⋅2⋅2⋅2⋅3 − − − − − − − − − − √ =2⋅23 √

OpenStudy (anonymous):

(2 rad.8)(3 rad.6)+19rad.48 6 rad.48+19 rad.48 25 rad.48 25 rad.16*3 25*4 rad.3 100 rad.3

OpenStudy (anonymous):

did you get it

OpenStudy (phi):

btw, when you said \[ 2 \sqrt{2}\sqrt{4} and 3 \sqrt{2}\sqrt{3} + 19 \sqrt{2}\sqrt{14} \] that should be \[ 2 \sqrt{2}\sqrt{4} \cdot 3 \sqrt{2}\sqrt{3} + 19 \sqrt{2}\sqrt{24} \] you could continue by saying \[ 2 \cdot 3 \sqrt{2}\cdot \sqrt{2}\cdot \sqrt{4}\sqrt{3} + 19 \sqrt{2}\sqrt{24} \] and the left part is 6 * 2 * 2 * sqrt(3) you would have to simplify the right part (the sqrt(24) ) to finish

OpenStudy (anonymous):

Rathanreddy I need you to draw the last 3 parts of your answer please

OpenStudy (anonymous):

Phi I am confused with the